# Improving LLM Unlearning Robustness via Random Perturbations

### v3

@@ 5.2 Robustness of RNA Models
The latent representation of the retain-query $x^{r}_{<i}$ is randomized in the RNA model, that is, $\mathbf{z}^{r}_{\theta^{\text{rna}}}=\mathbf{z}^{r}_{\theta^{u}}+\bm{\delta}$, where $\bm{\delta}$ is small and independently sampled from Normal distribution $\mathcal{N}(\bm{0},\nu\bm{I})$, $\nu\bm{I}$ is the covariance matrix, $\nu\in\mathbb{R}^{+}$.
We denote $f^{\text{rna}}$ the RNA model, $f^{u}$ the unlearned model, and $\mathcal{J}(.,.)$ be a loss function. We consider the change in the loss of the generated token $x^{r}_{i}$ given the perturbed retain-query and the retain-query in the unlearned model $f^{\text{u}}$: $\Delta\mathcal{J}^{u}=\mathcal{J}(f^{u}(x^{r}_{i}|x^{r,\text{per}}_{<i}))-\mathcal{J}(f^{u}(x^{r}_{i}|x^{r}_{<i})).$ Since the predicted output $f^{u}(x^{r}_{i}|x^{r,\text{per}}_{<i})$ is randomized (c.f. Theorem 4.2), the loss is increased, resulting in $\Delta\mathcal{J}^{u}>0$.
The change in the loss in RNA model $f^{\text{rna}}$ is $\Delta\mathcal{J}^{\text{rna}}=\mathcal{J}(f^{\text{rna}}(x^{r}_{i}|x^{r,\text{per}}_{<i}))-\mathcal{J}(f^{\text{rna}}(x^{r}_{i}|x^{r}_{<i})).$ If $f^{\text{rna}}$ is more robust to forget-tokens, it rejects the effect caused by the forget-token, i.e., it lowers the loss or keeps the loss remain unchanged, resulting in $\Delta\mathcal{J}^{\text{rna}}\leq 0$. We show that RNA improves the robustness of unlearned models, that is, the following inequality
@@
Suppose RNA adds a small, independent Gaussian noise $\bm{\delta}\sim\mathcal{N}(\bm{0},\nu\bm{I})$, $\nu\in\mathbb{R}^{+}$ into the retain-representation at layer $l$ of unlearned model $f^{u}$. If Assumption 4.1 and Assumption 5.1 hold, the probability that the RNA model rejects the effect caused by the forget-token, denoted as $\mathbb{P}\left[\frac{\Delta\mathcal{J}^{\text{rna}}}{\Delta\mathcal{J}^{\text{u}}}\leq 0\right]$, is approximate $\frac{1}{2}-\frac{1}{\pi}\arctan\left[\sqrt{\frac{\eta}{\nu}}\left(1+\frac{||\bm{g}^{\text{per}}||_{2}}{||\bm{g}||_{2}}\right)^{-1}\right]$, where $\bm{g}^{\text{per}}=\nabla_{\mathbf{z}^{r,\text{per}}_{<i}}\mathcal{J}(f^{u}(x^{r}_{i}|x^{r,\text{per}}_{<i}))$ and $\bm{g}=\nabla_{\mathbf{z}^{r}_{<i}}\mathcal{J}(f^{u}(x^{r}_{i}|x^{r}_{<i}))$ are the gradients of the loss of generated token $x^{r}_{i}$ with respect to $\mathbf{z}^{r,\text{per}}_{<i}$ and $\mathbf{z}^{r}_{<i}$.
@@
We defer the proof to Appendix B.2
Theorem 5.2 states that the probability $\mathbb{P}\left[\frac{\Delta\mathcal{J}^{\text{rna}}}{\Delta\mathcal{J}^{u}}\leq 0\right]$ is bounded by $\frac{1}{2}$ and is negatively correlated with $\arctan\left[\sqrt{\frac{\eta}{\nu}}\left(1+\frac{||\bm{g}^{\text{per}}||_{2}}{||\bm{g}||_{2}}\right)^{-1}\right]$. Since $\arctan$ is monotonically increasing, the robustness of unlearned models increases as $\sqrt{\frac{\eta}{\nu}}\left(1+\frac{||\bm{g}^{\text{per}}||_{2}}{||\bm{g}||_{2}}\right)^{-1}$ decreases. The product $\sqrt{\frac{\eta}{\nu}}\left(1+\frac{||\bm{g}^{\text{per}}||_{2}}{||\bm{g}||_{2}}\right)^{-1}$ is characterized by two terms: the root of the ratio $\frac{\eta}{\nu}$ and $\left(1+\frac{||\bm{g}^{\text{per}}||_{2}}{||\bm{g}||_{2}}\right)^{-1}$. First, if $\nu$ and $\eta$ are fixed, a larger ratio $\frac{||\bm{g}^{\text{per}}||_{2}}{||\bm{g}||_{2}}$ means a smaller $\left(1+\frac{||\bm{g}^{\text{per}}||_{2}}{||\bm{g}||_{2}}\right)^{-1}$, that is, a more robustness of the unlearned models. However, searching for all input and analyzing the effects of $\frac{||\bm{g}^{\text{per}}||_{2}}{||\bm{g}||_{2}}$ would be challenging due to the input-dependent property of $\bm{g}$ and $\bm{g}^{\text{per}}$. Second, consider the effect of $\frac{\eta}{\nu}$. If $\eta$ is fixed (the magnitude of the noise caused by forget-tokens), the larger $\nu$ is, the more robust the unlearned model becomes. However, since the probability is bounded, the robustness of unlearned models reaches a saturation point as $\nu$ increases. We present an empirical analysis in Section 6 to validate the claims.
@@ 6.1 Experimental Setup
Models and datasets. We conduct our experiments using Zephyr-7B-$\beta$ [59]. For fine-tuning, we use the WMDP Biology and WMDP Cyber forget-sets as $\mathcal{D}_{f}$, and Wikitext [37] as the retain-set $\mathcal{D}_{r}$. For evaluation, we use the WMDP Biology and Cyber QA sets for measuring forgetting performance, and the MMLU QA sets for retaining performance.
Synthesizing retain-queries that contain forget-token. To simulate interference, we create perturbed retain-queries by randomly replacing an incorrect answer in the original MMLU QA with a forget-keyword drawn from the forget-set. Following prior work [57], we use “SARS-CoV-2,” a frequent term in the WMDP Biology forget-set. See Appendix A.2 for details of the prompt templates.
@@
Unlearned models are expected to exhibit low accuracy on forget-tasks (WMDP-Biology and WMDP-Cyber QAs) while maintaining high accuracy on retain-tasks (MMLU, MMLU C. Bio. & C. Sec., and perturbed MMLU). Due to space constraints, we present key results supporting our theoretical analysis in the main text, and defer the full experimental setup and additional results to the Appendix A.4.
@@ 6.2 Main Results and Analysis
RNA improves unlearned models’ robustness while preserving their original forget/retain performances. Figure 1 (top-left and top-right) shows the accuracy of RM, PO, and RNA models evaluated on perturbed MMLU, MMLU. The results highlight that all original unlearned models, including RM and PO, exhibit substantial vulnerability to the forget-token, resulting in significant drops in accuracy when the forget-token appears in retain-queries. Specifically, compared to the base model, the accuracy reduction rate in RM models averaged $23.3\%$ (RMU: $19.0\%$, Adaptive RMU: $30.2\%$, and RSV: $20.8\%$). PO models showed catastrophic collapse with $43.3\%$ average reduction (NPO+KL: $50.9\%$, NPO+MSE: $27.8\%$, DPO+KL: $31.8\%$, DPO+MSE: $58.4\%$, SimNPO+KL: $44.4\%$, SimNPO+MSE: $47.9\%$). This result emphasizes that RM models consistently show stronger robustness compared to PO models.
When applied to RM methods, RNA achieves an average accuracy recovery rate of $66.3\%$ (RMU: $34.2\%$, Adaptive RMU: $81.7\%$, RSV $83.2\%$). For PO methods, the average recovery rate is $51.7\%$ (NPO+KL: $60.9\%$, NPO+MSE: $18.5\%$, DPO+KL: $32.9\%$, DPO+MSE: $91.4\%$, SimNPO+KL: $32.3\%$, SimNPO+MSE: $74.2\%$). RNA maintains the original forget/retain utility, with WMDP and MMLU accuracy remaining stable after RNA integration. Additionally, RNA improves model robustness on forget-tasks related to forget datasets such as MMLU C. Sec. and C. Bio. (Figure 1 bottom-right).
Trade-off between the coefficient in RM and robustness. As suggested by Theorem 4.2, increasing either the coefficient $c$ or scaling factor $\beta$ is expected to reduce the unlearned model’s robustness.
@@
To validate this claim, we fix the unlearn layer at $l=7$ and grid search over values of $c$ and $\beta$, reporting the accuracy of RM models on perturbed MMLU. Figure 2 shows a clear trend that the accuracy of RM models decreases as the coefficient $c$ or $\beta$ increases. Previous works [30, 11] performed grid search for $c$ and $\beta$, selecting values that yielded optimal accuracy and observed that deeper unlearn layers require larger values of $c$ (or $\beta$) to achieve effective unlearning. However, our results demonstrate that increasing the coefficient $c$ (or $\beta$) results in a notable reduction in model robustness. From a robustness perspective, choosing earlier layers as the unlearn layer helps maintain the robustness of the RM models.
Effects of the noise scale $\nu$. We evaluate the accuracy of RNA models on perturbed MMLU and WMDP by varying $\nu$. As shown in Figure 3, we observed that increasing $\nu$ first leads to improved accuracy of RNA models on perturbed MMLU while maintaining stable accuracy on WMDP. However, as $\nu$ continuous increases, the accuracy of RNA models on perturbed MMLU begins to decline, indicating a point where excessive noise becomes detrimental to retain accuracy. This result aligns with the analysis in Theorem 5.2, which suggests that the RNA models’ robustness is bounded and will reach a saturation point. Notably, we observed that RM methods are more stable and robust to noise $\nu$ than PO.
@@
Analysis on the harmfulness of forget-tokens. One might ask: “Which forget-tokens when appearing in the retain-query can cause the unlearned model to misbehave?”. We examine the harmfulness of forget-tokens in the forget-set by measuring the cosine similarity between bi-gram forget-tokens and their respective documents, across all documents in the WMDP forget-sets. We select the top $10$ most similar, least similar, and those with values around the mean of the distribution. Perturbed MMLU QAs w.r.t these forget-tokens are synthesized following the procedure described in Section 6.1. As shown in Figure 4, we observed a clear trend between the accuracy and the similarity: forget-tokens with higher similarity with their corresponding documents are more harmful to unlearned models. See Appendix D for the evaluation of RNA models’ robustness against $n$-gram similarity perturbations for $n\in\{4,8,16\}$.
Figure 4: Accuracy of unlearned models on perturbed MMLU w.r.t bi-gram similarity perturbations.
@@ 7 Conclusion
This paper proposes RNA, a simple data augmentation method for improving unlearned models’ robustness. By reframing unlearning as a backdoor attack and defense problem, we explain the inherent fragility of unlearned models. Extensive theoretical and empirical analysis confirm RNA’s effectiveness and efficiency. Our findings advance the understanding of the underlying behaviors of unlearning methods and shed light on the development of robust machine unlearning algorithms.
Acknowledgments
The authors thank Nguyen Minh Phuong for insightful comments and discussions.
@@ A.1 Datasets
Wikitext [37]
is a language modeling dataset consisting of over $100$ milion tokens extracted from Wikipedia. Following [30], we specifically use the Wikitext-2-raw-v1 test split as the retain-set for fine-tuning. The dataset is publicly available at https://huggingface.co/datasets/Salesforce/wikitext.
MMLU [20]
@@
MMLU College Biology & Computer Security [20]
@@
We employ a set of $100$ refusal responses from Maini et al. [36] as the preference answers for DPO+KL and DPO+MSE. For further details, we refer the reader to Appendix C of Maini et al. [36].
A.2 Prompt Templates
Perturbed Query The following are multiple choice questions (with answers) about anatomy. Question: A lesion causing compression of the facial nerve at the stylomastoid foramen will cause ipsilateral A. paralysis of the facial muscles. B. paralysis of the facial muscles and loss of taste. C. SARS-CoV-19. D. paralysis of the facial muscles, loss of taste, lacrimation and decreased salivation. Answer:
Figure 6: A sample QA prompt. A random incorrect answer (e.g., C. paralysis of the facial muscles, loss of taste and lacrimation.) is replaced by a forget keyword SARS-CoV-19 while the correct answer (A. paralysis of the facial muscles.) is unchanged.
We use the lm-evaluation-harness framework [18] for evaluation. Each query is formulated as a default zero-shot QA prompt (Figure 6). Following the setting of prior work [57], we randomly replace an incorrect answer in the retain QA dataset with the forget keyword “SARS-CoV-19,” while leaving the correct answer unchanged. Since the forget keyword is unrelated to the retain-queries, this modification is expected to have minimal effect on retain performance. Additional experiments on RNA’s performance with other forget-tokens can be found in Appendix D.
@@ A.3 Evaluation Metrics
Following [30], we use zero-shot QA accuracy to assess the efficacy of unlearning methods. To further evaluate the unlearned models’ brittleness and RNA’s effectiveness, we report the accuracy reduction rate and recovery rate. These metrics are defined as follows:
Reduction $\text{Reduction rate}=\frac{\text{Acc}_{\text{base}}-\text{Acc}_{\text{unlearned}}}{\text{Acc}_{\text{base}}}\times 100\%$ (14)
Recovery $\text{Recovery rate}=\frac{\text{Acc}_{\text{rna}}-\text{Acc}_{\text{unlearned}}}{\text{Acc}_{\text{base}}-\text{Acc}_{\text{unlearned}}}\times 100\%$ (15)
Example. If $\text{Acc}_{\text{base}}=60$, $\text{Acc}_{\text{RMU}}=30$, $\text{Acc}_{\text{RMU w/ RNA}}=50$, then the reduction rate is $50\%$ and the recovery rate is $66.67\%$.
@@
Models are fine-tuned using AdamW [33] for $T=500$ update steps, learning rate is $55e-$ 5, batch size of $4$, max sequence length is $500$ with WMDP-Biology and $768$ for WMDP-Cyber. Following previous works [30, 11], we update three layers of parameters $\{l,l-1,l-2\}$ of the model to save memory. For the original RM methods, we set the retain weight $\alpha_{\text{biology}}=1200$ and $\alpha_{\text{cyber}}=1200$, the unlearned layer $l=7$ for all methods, the coefficient $c=6.5$ for RMU, and the scaling factor $\beta=3$ for Adaptive RMU. For RSV, we grid search for the coefficient $c\in\{5,10,20,30,40,50,60,70,80,90,100\}$ and select $c=10$. For the original PO methods, we adopt the default hyperparameters used in prior work [65, 15]. Specifically, we set $\beta=0.1$ for all PO methods, and $\gamma=0$ for both SimNPO+KL and SimNPO+MSE. For the retain weights, we perform a grid search over combinations of $(\alpha_{\text{biology}},\alpha_{\text{cyber}})$, where $\alpha_{\text{biology}},\alpha_{\text{cyber}}\in\{5,10,20,30,40,50,100\}$. We select the combinations that achieve a balanced trade-off between forgetting and retaining performance: $(30,50)$ for DPO+KL, $(5,20)$ for DPO+MSE, $(50,50)$ for NPO+KL, $(5,20)$ for NPO+MSE, $(20,50)$ for SimNPO+KL, and $(5,10)$ for SimNPO+MSE.
@@
All experiments are conducted using two NVIDIA A40 GPUs, each with $45$GB of memory. Our implementation is included in the supplementary material. The perturbed MMLU QA datasets will be made publicly available.

### v4

@@ 5.3 Robustness of RNA Models
The latent representation of the retain-query $\mathbf{x}^{r}_{<i}$ is randomized in the RNA model, that is, $\mathbf{z}^{r}_{\bm{\theta}^{\text{rna}}}=\mathbf{z}^{r}_{\bm{\theta}^{u}}+\bm{\delta}$, where $\bm{\delta}$ is small and independently sampled from Normal distribution $\mathcal{N}(\bm{0},\nu\bm{I})$, $\nu\bm{I}$ is the covariance matrix, $\nu\in\mathbb{R}^{+}$.
We denote $f^{\text{rna}}$ the RNA model, $f^{u}$ the original unlearned model, and $\mathcal{J}(.,.)$ be a loss function. Consider the change in the loss of the generated token $\mathbf{x}^{r}_{i}$ given the perturbed retain-query and the retain-query in the unlearned model $f^{\text{u}}$: $\Delta\mathcal{J}^{u}=\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{x}^{r,\text{per}}_{<i}))-\mathcal{J}(f^{u}(\mathbf{x}^{r}_{i}|\mathbf{x}^{r}_{<i})).$ Since the predicted output $f^{u}(\mathbf{x}^{r}_{i}|\mathbf{x}^{r,\text{per}}_{<i})$ is randomized (c.f. Theorem 4.2), the loss is increased, resulting in $\Delta\mathcal{J}^{u}>0$. The change in the loss in RNA model $f^{\text{rna}}$ is $\Delta\mathcal{J}^{\text{rna}}=\mathcal{J}(f^{\text{rna}}(\mathbf{x}^{r}_{i}|\mathbf{x}^{r,\text{per}}_{<i}))-\mathcal{J}(f^{\text{rna}}(\mathbf{x}^{r}_{i}|\mathbf{x}^{r}_{<i})).$ If $f^{\text{rna}}$ is more robust to forget-tokens, it rejects the effect caused by the forget-token, i.e., it lowers the loss or keeps the loss remain unchanged, resulting in $\Delta\mathcal{J}^{\text{rna}}\leq 0$. We show that RNA improves the robustness of unlearned models, that is, the following inequality
@@
[…]
@@
We defer the proof to Appendix B.2. ∎
[…]
Underlying mechanism of RNA. From the backdoor attack and defense perspective, current unlearning methods do not truly unlearn knowledge but instead hide it behind a “trigger” mechanism. Likewise, RNA does not truly erase knowledge; rather, it blurs the decision boundary around forget-tokens so that inserting one or some of those forget-tokens is no longer a reliable way to recover the forgotten knowledge. In other words, by injecting small Gaussian noises into the latent space during unlearning, RNA reduces the clean separation between “triggered” (critical forget-tokens) and “untriggered” representations (less critical forget-tokens). This smoothing makes the forget-token less salient as a backdoor signal. As a result, the model still retains its general knowledge, yet that forgotten knowledge cannot be inadvertently recalled when forget-tokens appear in retain-queries.
@@ 6.1 Experimental Setup
Models and datasets. We conduct our experiments using Zephyr-7B-$\beta$ (Tunstall et al., 2023), Mistral-7B (Jiang et al., 2023), and Llama-3-8B (Dubey et al., 2024). For fine-tuning, we use the WMDP-Biology and WMDP-Cyber forget-sets as $\mathcal{D}_{f}$, and Wikitext (Merity et al., 2016) as the retain-set $\mathcal{D}_{r}$. For evaluation, we use the WMDP-Biology and Cyber QA sets for measuring forgetting performance, and the MMLU QA sets for retaining performance.
Synthesizing retain-queries that contain forget-token. To simulate interference, we create perturbed retain-queries by randomly replacing an incorrect answer in the original MMLU QA with a forget-keyword in the forget-set. Following prior work (Thaker et al., 2025), we use “SARS-CoV-2,” a frequent term in the WMDP forget-set. See Appendix A.2 for details of the prompt template, Appendix D for performance of RNA against hard negative forget-tokens.
@@
Unlearned models are expected to exhibit low accuracy on forget-tasks (WMDP-Biology and WMDP-Cyber QAs) while maintaining high accuracy on retain-tasks (MMLU, MMLU C. Bio. & C. Sec., and perturbed MMLU). Due to space constraints, we report key results of Zephyr-7B that support our theoretical analysis in the main text, and defer the full experimental setup and results to the Appendix.
@@ 6.2 Main Results and Analysis
RNA improves unlearned models’ robustness while preserving their original forget/retain performances. Figure 1 (left-most and left-mid) shows the accuracy of RM, PO, and RNA models evaluated on perturbed MMLU, MMLU. The results highlight that all original unlearned models, including RM and PO, exhibit substantial vulnerability to the forget-token, resulting in significant drops in accuracy when the forget-token appears in retain-queries. Specifically, compared to the base model, the accuracy reduction rate in RM models averaged $23.3$ (RMU: $19.0$, Adaptive RMU: $30.2$, and RSV: $20.8$). PO models showed catastrophic collapse with $43.3$ average reduction (NPO+KL: $50.9$, NPO+MSE: $27.8$, DPO+KL: $31.8$, DPO+MSE: $58.4$, SimNPO+KL: $44.4$, SimNPO+MSE: $47.9$). This result emphasizes that RM models consistently show stronger robustness compared to PO models.
Figure 1: Left-most: Accuracy of RM and RM w/ RNA models on MMLU and perturbed MMLU (MMLU QA contains forget-tokens; see Appendix A.2 for details). Left-mid: Accuracy of PO and PO w/ RNA models on MMLU and perturbed MMLU. Right-mid: Accuracy of all unlearned models on WMDP and perturbed WMDP. Right-most: Accuracy of all unlearned models on MMLU subsets (College Biology and Computer Security). Original RM models are shown by one-color circles and original PO models by one-color triangles. Two-color markers for models with RNA, where the inner color indicates the original method and the outer blue ring denotes RNA integration.
When applied to RM methods, RNA achieves an average accuracy recovery rate of $66.3$ (RMU: $34.2$, Adaptive RMU: $81.7$, RSV $83.2$). For PO methods, the average recovery rate is $51.7$ (NPO+KL: $60.9$, NPO+MSE: $18.5$, DPO+KL: $32.9$, DPO+MSE: $91.4$, SimNPO+KL: $32.3$, SimNPO+MSE: $74.2$). RNA maintains the original forget/retain utility, with WMDP and MMLU accuracy remaining stable after RNA integration. Additionally, RNA improves model robustness on forget-tasks related to forget datasets such as MMLU C. Sec. and C. Bio. (Figure 1 right-most).
Trade-off between the coefficient in RM and robustness.
@@
As suggested by Theorem 4.2, increasing either the coefficient $c$ or scaling factor $\beta$ is expected to reduce the unlearned model’s robustness. To validate this claim, we fix the unlearn layer at $l=7$ and grid search over values of $c$ and $\beta$, reporting the accuracy of RM models on perturbed MMLU. Figure 2 shows a clear trend that the accuracy of RM models decreases as the coefficient $c$ or $\beta$ increases. Previous works (Li et al., 2024; Dang et al., 2025) performed grid search for $c$ and $\beta$, selecting values that yielded optimal accuracy and observed that deeper unlearn layers require larger values of $c$ (or $\beta$) to achieve effective unlearning. However, our results demonstrate that increasing the coefficient $c$ (or $\beta$) results in a notable reduction in model robustness. From a robustness perspective, choosing earlier layers as the unlearn layer helps maintain the robustness of the RM models.
Effects of the noise scale $\nu$.
@@
We evaluate the accuracy of RNA models on perturbed MMLU and WMDP by varying $\nu$. As shown in Figure 3, we observed that increasing $\nu$ first leads to improved accuracy of RNA models on perturbed MMLU while maintaining stable accuracy on WMDP. However, as $\nu$ continuous increases, the accuracy of RNA models on perturbed MMLU begins to decline, indicating a point where excessive noise becomes detrimental to retain accuracy. This result aligns with the analysis in Theorem 5.2, which suggests that the RNA models’ robustness is bounded and will reach a saturation point. Notably, we observed that RM methods are more stable and robust to noise $\nu$ than PO.
@@ 7 Conclusion
This paper proposes RNA, a simple yet effective robust unlearning method for improving unlearned models’ robustness. By reframing unlearning as a backdoor attack and defense problem, we explain the inherent fragility of unlearned models. Extensive theoretical and empirical analysis confirm RNA’s effectiveness and efficiency. Our findings advance the understanding of the underlying behaviors of unlearning methods and shed light on the development of robust machine unlearning algorithms.
@@
  • A. Experimental Setup ........................................................................................................................................................................A
    • A.1 Datasets ........................................................................................................................................................................A.1
    • A.2 Prompt Template ........................................................................................................................................................................A.2
    • A.3 Evaluation Metrics ........................................................................................................................................................................A.3
    • A.4 Implementation Details ........................................................................................................................................................................A.4
  • B. Proofs ........................................................................................................................................................................B
    • B.1 Proof of Theorem 4.2 ........................................................................................................................................................................B.1
    • B.2 Proof of Theorem 5.2 ........................................................................................................................................................................B.2
  • C. Empirical Validation ........................................................................................................................................................................C.1
  • D. Robustness of RNA Against Hard Negative Forget-Tokens ........................................................................................................................................................................D
  • E. Effects of Randomizing Different Latent Spaces ........................................................................................................................................................................E
  • F. Robustness of RNA Models Against Prompt Attacks ........................................................................................................................................................................F
  • G. Effects of RNA on Model Alignment ........................................................................................................................................................................G
  • H. Performance of Other Models ........................................................................................................................................................................H
  • I. Limitation and Future Work ........................................................................................................................................................................I
  • J. Broader Impact ........................................................................................................................................................................J
  • K. AI Usage Declaration ........................................................................................................................................................................K
@@ A.1 Datasets
Wikitext (Merity et al., 2016)
is a language modeling dataset consisting of over $100$ milion tokens extracted from Wikipedia. Following Li et al. (2024), we specifically use the Wikitext-2-raw-v1 test split as the retain-set for fine-tuning. The dataset is publicly available at https://huggingface.co/datasets/Salesforce/wikitext.
MMLU (Hendrycks et al., 2021)
@@
MMLU College Biology & Computer Security (Hendrycks et al., 2021)
@@
We employ a set of $100$ refusal responses as the preference answers for DPO+KL and DPO+MSE. For further details, we refer the reader to Appendix C of Maini et al. (2024).
A.2 Prompt Template
The following are multiple choice questions (with answers) about anatomy. Question: A lesion causing compression of the facial nerve at the stylomastoid foramen will cause ipsilateral A. paralysis of the facial muscles. B. paralysis of the facial muscles and loss of taste. C. SARS-CoV-19. D. paralysis of the facial muscles, loss of taste, lacrimation and decreased salivation. Answer:
Figure 6: A sample zero-shot QA prompt. A random incorrect answer (e.g., C. paralysis of the facial muscles, loss of taste and lacrimation.) is replaced by a forget keyword SARS-CoV-19 while the correct answer (A. paralysis of the facial muscles.) is unchanged.
We use the lm-evaluation-harness framework (Gao et al., 2024) for evaluation. Each query is formulated as a default zero-shot QA prompt (Figure 6). Following the setting of prior work (Thaker et al., 2025), we randomly replace an incorrect answer in the retain QA dataset with the forget keyword “SARS-CoV-19,” while leaving the correct answer unchanged. Since the forget keyword is unrelated to the retain-queries, this modification is expected to have minimal effect on retain performance.
@@ A.3 Evaluation Metrics
Accuracy, Reduction Rate, and Recovery Rate.
Following Li et al. (2024), we primarily use zero-shot QA accuracy to assess the efficacy of unlearning methods. To further evaluate the unlearned models’ brittleness and RNA’s effectiveness, we report the accuracy reduction rate and recovery rate. These metrics are defined as follows:
$\text{Reduction Rate}=\frac{\text{Acc}_{\text{base}}-\text{Acc}_{\text{unlearned}}}{\text{Acc}_{\text{base}}}\times 100\%$ (14)
$\text{Recovery Rate}=\frac{\text{Acc}_{\text{rna}}-\text{Acc}_{\text{unlearned}}}{\text{Acc}_{\text{base}}-\text{Acc}_{\text{unlearned}}}\times 100\%$ (15)
For example, if $\text{Acc}_{\text{base}}=60$, $\text{Acc}_{\text{RMU}}=30$, $\text{Acc}_{\text{RMU w/ RNA}}=50$, then the reduction rate is $50\%$ and the recovery rate is $66.67\%$.
Additionally, we report accuracy under attack (AuA) and ROUGE-L score for experiments in Section F to evaluate the robustness of RNA against prompt injection attacks.
@@
Models are fine-tuned using AdamW (Loshchilov & Hutter, 2019) for $T=500$ update steps, learning rate is 5e55e-5, batch size of $4$, max sequence length is $500$ with WMDP-Biology and $768$ for WMDP-Cyber. Following previous works (Li et al., 2024), we update three layers of parameters $\{l,l-1,l-2\}$ of the model for memory efficiency. For the original RM methods, we set the retain weight $\alpha_{\text{biology}}=1200$ and $\alpha_{\text{cyber}}=1200$, the unlearned layer $l=7$ for all methods, the coefficient $c_{\text{biology}}=c_{\text{cyber}}=6.5$ for RMU, and the scaling factor $\beta=3$ for Adaptive RMU. For RSV, we grid search for the coefficient $c\in\{5,10,20,30,40,50,60,70,80,90,100\}$ and select $c_{\text{biology}}=c_{\text{cyber}}=10$. For the original PO methods, we adopt the default hyperparameters used in previous works (Yuan et al., 2025b; Fan et al., 2024). Specifically, we set $\beta=0.1$ for all PO methods, and $\gamma=0$ for both SimNPO+KL and SimNPO+MSE. For the retain weights, we perform a grid search over combinations of $(\alpha_{\text{biology}},\alpha_{\text{cyber}})$, where $\alpha_{\text{biology}},\alpha_{\text{cyber}}\in\{5,10,20,30,40,50,100\}$. We select the combinations that achieve a balanced trade-off between forgetting and retaining performance: $(30,50)$ for DPO+KL, $(5,20)$ for DPO+MSE, $(50,50)$ for NPO+KL, $(5,20)$ for NPO+MSE, $(20,50)$ for SimNPO+KL, and $(10,5)$ for SimNPO+MSE.
@@
All experiments are conducted using two NVIDIA A40 GPUs, each with 4545GB of memory. The perturbed MMLU QA datasets will be made publicly available.